Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, November 17, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 17, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/17/98 (Tuesday)

Segment 1 (Emidio to ANF). Report for 17 November, 1998. Work continued at the bore pits at the Cuddy creek tributary on Peace Valley Road, ~EMP 25.45. At the skip on Lebec Road, one hoe excavated at both ends of the skip, the hoe to the south progressed ~ 40 feet from EMP 21.75. The northern excavation move southward another 60 feet, approximately. Work continued on the Tejon Bench, at EMP 18.51, excavation at EMP 18.15 and more crews working to the north. On Grapevine Road, the excavation ended on the road in front of Grapevine Station, ~EMP 15.58 The bell holes were excavated to immediately south of that location. The bending machine was set up at EMP 16.00. with pipe lowering and welding to the south of that. Work continued at the bore sites. Recontouring work along Access Road B continued. Work along Highway 166 consisted of trenching and shoring east of ~EMP 1.0. Backfill worked near EMP .33.

Segment 2 (ANF to Sylmar) Report for 11/17/98: 2A (Copco Rd to Templin Hwy): The drying 'pig' run that was started yesterday from the Copco Rd cutout at EMP 33.35 reached the Marple Cyn cutout, EMP 55.46 in Segment 2C. 2C (Templin Hwy to Castaic): A meeting was held to discuss the status of the Project within Angeles National Forest land in Segments 2A and 2C. Present were representatives of Angeles National Forest, PPSI, LWReed, ARB, and Aspen Environmental Group. The ROW over the Templin Hwy testing cutout site, which has just been recontoured, was surveyed for replacement of the dirt road present before the project. At the Marple Cyn cutout, EMP 55.46, there was considerable quantity of water pushed in front of the drying 'pig'. This water was discharged into Marple Creek, producing temporary increase in flow and turbidity. Recontouring took place near EMP 55.50. At Lake Hughes block valve, EMP 56.62, work began on the anchor forms. 2B (Castaic to Sylmar): Trenching took place at Romero Cyn, EMP 59.16, for electrical attachments. Boring continued at Feedmill Rd. on the Santa Clara River directional drill. Crews also pumped drilling mud out of a frac-out in the river bed, and placed berms around that and another frac-out site. At Calgrove, x-ray and coating were completed at the tie-in. Fabricated pipe that had been welded at the Calgrove work area was hauled off to the south end of Segment 2B. Cold planing of the roadway south of Weldon Cyn was completed south to EMP 72.23. Trenching proceeded from the south northward to EMP 72.35. Pipe was strung in the trench to 72.50, and tied-in to Sierra Hwy. Backfilling with aggregate base reached EMP 72.9. General clean-up including removal of shoring jacks occurred south of there.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 17 November 1998: 3A: Backfill crews continued their work on either side of the bridge crossing of San Fernando Road. The crew also began work on either side of the bridge crossing at Hollywood Way, including fiber optic conduit placement. Welding, x-ray, and coating all spent the day working on the tie in of the Van Nuys bore. Minor work was done at Tujunga Wash, in anticipation of placing pipe for this bridge crossing. Several brackets were hung on the railroad bridge for Pacoima Wash crossing. Fiber optic crews spent the day in Glendale, blowing rope into the four conduit. The bore crew at Tuxford Street continued setting up. A crew worked late digging assist pits in Burbank and Glendale. 3B: No work occurred. 3C: The tie in to the Hubbard bore continued today in San Fernando. Backfill crews poured slurry over the completed pipe in San Fernando Road, just past Sepulveda Boulevard. Work on the aqueduct crossing progressed, including work on the high speed channel cover on the west side of the road. Crew dug out between the railroad bore and the aqueduct connection site.

Segment 4B(El Segundo to Lynwood) 11-17-98: Construction activities continued at El Segundo Station ESEMP 11.00 . The crew at Western ESEMP 5.56 was backfilling. The crew inside Lynwood station were working on the south wall. The paving crew was working at 108 Th St. ESEMP 1.11, and at Alameda and Fernwood EMP 0.01.

Segment 5 (Lynwood to Long Beach) Report for 17 November 1998: Directional drill at Highway 91/Compton creek continued with 36-inch reamer. Boring crew at Wilmington station pushed though two joints of pipe and repaired pipe at pot-hole.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 4 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page